mingdale.open.suny s m it take saunch/sp?course a sessment d# 25064 1 &course, d 18242 1 &content id 576217 1 &step nu QUESTION 28 Speed rato af a chain rive s selected as 3.1, If he number of teeth on diver sprocket is 17, what is the number of teeth on the drive n sprocket ? 0 17 O None of the above QUESTION 29 The service factor for a V-Belt system where a bucket olevator is driven by an AC motor with high torque for 8 hours per day is 12 1.3 0 14 0 15 QUESTION 30 e service factor for a chain drive used for a heavy conveyor driven by an eiectri

media%2Fc6e%2Fc6e7b711-f619-4bba-9f81-e5
Answer #1

28 )

Speed ratio = 3 :1 = (driven sprocket teeth) / (driver sprocket teeth) = T2 / T1 = T2 / 17

Therefore, T2 = 17 * 3 = 51

Answer is 51

29 ) From the design data book, for high torque AC Motors and used in bucket elevators for 8 hours a day

The service factor for V- belt is 1.5

Answer is 1.5

30)  From the design data book, for heavy conveyor driven by electric motor,

The service factor is 1.5

answer is 1.5